Lindsay​ Electronics, a small manufacturer of electronic research​ equipment, has approximately

7 comma 0007,000

items in its inventory and has hired Joan​ Blasco-Paul to manage its inventory. Joan has determined that

1111​%

of the items in inventory are A​ items,

3232​%

are B​ items, and

5757​%

are C items. She would like to set up a system in which all A items are counted monthly​ (every

2020

working​ days), all B items are counted quarterly​ (every

5959

working​ days), and all C items are counted semiannually​ (every

120120

working​ days). How many items need to be counted each​ day?

The total number of items that need to be counted each day is

nothing

items ​(round your response to the nearest whole​ number).
Answer #1

Note: Please rate the answer (upvote) and do ask if any doubt is there.

Answer:

  • A items are 0.11* 7000 = 770
  • B items are 0.32 * 7000 = 2240
  • C Items are 0.57 * 7000 = 3990
  • Units counted per day to be counted:
    • A: Monthly: 770/20
    • B: quarterly: 2240/60
    • C: semi-annually: 3990/120
  • Units to be counted everyday= 770/20 + 2240/60 + 3990/120 = 109.08 or 109 items (Answer)

Note: Please rate the answer (upvote) and do ask if any doubt is there.